Startup Night 4 - 20140902 UT

Observers: JHill (LBTO), OKuhn (Tucson)
Telescope Operator: SAllanson (LBTO)
AO Support: JCGuerra
Telescope Support: KNewton
Software Support: DSummers
Instrument Support: JPower (LBTO)
Engineering Support: DAshby

Summary

Pseudo-monocular mode MODS + short (15-s, 8-s and 4-s) LBCR observations were taken and this appeared to work OK. Further analysis needed to check for things like drift.

Took more superfoc sequences with LBCR because we again saw a focus and image quality gradient from top to bottom (this is what has typically been seen, on blue as well as on red --- images can be round at the top of chip 2 and elongated at the bottom). During the series of science exposures, John tweaked the Z value by 40 mic to get round images at the bottom and elongated ones at the top. dZ = +40 mic made a difference.

LBC-Red shutter was failing intermittently. A pop up would appear and the image would be lost. The images were very short: 15-sec, 8-sec and then 4-sec, and the OB was continuously being played by a python script, so this was not so noticeable. Although, around the time we started the 4-sec series, the shutter failed every time. It was not recovered for the rest of the night.

Obtained dual grating spectra of several spectrophotometric standards at a range of airmasses: 1.53, 1.45, 1.02, 1.4, 1.4, 2.1, 1.14, 1.285 (twilight sky may have been too bright for this one to be useful). Night was photometric.

Observed NGC 6341 (M92) field with ugriz photometry from Clem+2007. (this was the same field observed during MODS2 AGw commissioning).

Observed two more MODS science targets. Some observations:
  • sometimes a negative dX, ~0.2", is needed to tweak centering after modsAlign. So far during restart, I have not had to tweak centering by giving a postive dX value.
  • imcs seems slow to lock on the red side, although counts are fine. And the two times we were at airmass ~2, it would not lock within the timeout period and the script gave an error, though upon "R"etry, it would lock.
  • reference position for 5" slit should be checked - it seems more like 499/500 than 502.
  • collimation: guide star PSF changing and not round - most frequently has a sort of "arrowhead" shape. MODS images on-axis sometimes also not round. A good series to check would be the ones taken at the beginning of the night after the spectroscopic observation.

On the left edge of mods1r images, at Y=820, there is a dark shadow which extends to X~190. Will have to check whether it was there earlier.

Preparations

Jenny found a bad 120V adapter plug that was causing the LBC-Red Dataprobe issues.

Marco Stangalini sent us an updated version v0.9.6 of WRS based on last night's tests (it was not used this night).

The Plan for Tonight

The weather is looking great, so tonight's activities will be similar to last night. The telescope configuration will be MODS1 on the left and LBC-Red on the right in pseudomono parallel.

Olga and Barry will continue science verification and calibration activities with MODS1. John will work on LBC-Red. We'll test the MODS+LBC-Red transit observations in pseudomono mode. In the second half we'll test the new version of WRS.

Details

Open

01:45 Open to perfectly photometric sky; T=13 degC, D=-1.6 degC, wind light from W.

MODS1 on left, LBC-Red on right in pseudomono parallel

02:02 Red_SkyFlatScalings

LBC crashed

2014/09/02 02:02:38.508938 N R                   TCS IIF_GetRotatorTrajectory start
2014/09/02 02:02:38.508982 N R                   TCS IIF_GetRotatorTrajectory beginning at 56902.085162MJD and 346.000000 seconds long
2014/09/02 02:02:38.511281 N -                   issuing >tftp -m binary 192.168.38.101 -c put /lbccontrol/current/obs/lbc.140902.020238.ob lbc.140902.020238.ob >/dev/null 2>&1< upload command completed in 0.0 sec
2014/09/02 02:02:38.511323 S -                   file upload queue contains 0 file/s
2014/09/02 02:02:38.581797 N B                   Using TCS rotator trajectory [src/tcs/tcs.c:747]
2014/09/02 02:02:38.581831 N B                   TCSQueryRotatorTrajectory completed


2014/09/02 02:02:48.685664 < Log Service: session started >
2014/09/02 02:02:48.685828 < Log Service: RPC server started >
2014/09/02 02:02:48.686673 N -                   Starting LBC software version 2.0 - Aug 31 2014
2014/09/02 02:02:48.686711 N -                   clear ProfiBus control handles
2014/09/02 02:02:48.691362 N -                   local data initialize ...
2014/09/02 02:02:48.691446 N -                   local data initialize completed
2014/09/02 02:02:48.691460 N -                   accessing /lbccontrol/current/conf/lbc.conf configuration file ...
2014/09/02 02:02:48.691704 S -                   read ENABLE parameter >LBC< with value TRUE
2014/09/02 02:02:48.691730 S -                   read ENABLE parameter >LBC.UseTelescope< with value TRUE

Repeated the same crash at 02:10:52 -- See Issue 5285.

Also power.conf fails to reauthorize if only on one side.

02:15 DOFPIA, /REDONLY - sky is still too bright

02:20 acqMODS bd284211.acq (to do with execMODS -b DualGrating) still twilight
  • seeing not good, about 1.6".
  • on 0003, star to right of slit. offset by dx = -0.5" 0004 OK

02:23 Red_FlatScalings succeeds if I uncheck useblue.

02:27 DOFPIA, /FIRST also crashes because I didn't specify /REDONLY. See Issue 5286. These crashes made core files.

02:32 execMODS -b DualGrating bd284211.obs. Airmass 1.526, seeing 1.6"
  • guide camera images around 02:41 UT look like "frogs", according to John. Steve's clearing Active Optics. Meanwhile seeing has improved to about 1.2". AGw x= -72, y= -135.
  • repeating execMODS -b DualGrating bd284211 again (airmass ~1.45) after they cleared active optics and seeing improved. But collimation is not so great still - guide star sometimes looking like "arrowheads" 02:49 UT.

02:55 acqMODS PDS456k1.acq and sent PDS456preset.ob (but this was ignored - not necessary to send preset in psuedomono mode... I should have known...). LBC preset failed but took an image - LBC should have waited until it was "on-target", but it didn't.
  • LBC-Red shutter failure
  • 11 slit in, 12 field, collimation still a bit off 03:00. Centered at Y=630 instead of Y=590.

03:01:11 "Camera Shutter Hardware failure" (second one tonight)

03:09:06 lbcred image was fine with correct shutter telemetry


Computed Slit Alignment Offset:
   dX:  -4.135 arcsec
   dY: -17.302 arcsec


MODS Offset Command: 
   offsetxy -4.135 -17.302 rel

Send the offset to the Telescope (Y|N)?: Y

03:28 DOFPIA, /REDONLY

03:35:06 another camera shutter hardware failure.

2014/09/02 03:35:06.730604 W R CAMERA SHUTTER shutter position status returns >XPS:0, YPS:360003< [src/camera/camera.c:1247] 2014/09/02 03:35:06.730724 E R camera system encountered a problem: shutter hardware failure (10) [src/channel/channel.c:1215]

2014/09/02 03:41:39.737000 W R CAMERA CCDCTRL measured exposure time not available due to hardware timer failure, 8.000 sec estimated

03:47 +0.040 Z (seems correct when images are more elongated at bottom)

03:49 +0.040 Z

03:51 +0.040 Z

2014/09/02 03:52:20.869056 E R camera system encountered a problem: shutter hardware failure (10) [src/channel/channel.c:1215]

03:55 -0.040 Z

  • 13 looks good, but give -0.1" offset in X. 14 a little worse. 0.1" to the right again for 15.
  • execMODS PDS456_900sec.obs
    • seeing is 0.9-1"
    • LBC exposures are saturated on qso - reduce exptime from 15 to 8 seconds, in between which John ran dofpia
  • repeated lbc script: ./loop_observe_pds.py (in /home/lbto/supportscripts/SymbioticSupport/)
    • qso still saturated, will drop to 4 sec after for a final series later.
    • LBC images look elongated on axis and there is a definite gradient from top (round) to bottom (most elongated).
      • John giving offsets to Z of +40 mic focus, and again +40 mic.
  • UT 03:57 stopping loop_observe_pds.py script so we can run some superfoc seqeunces here. We want to confirm that the focus gradient is the same independent of PA, and try to measure it. Seeing is sub-arcsec.

04:01 DOFPIA, /REDONLY

04:04 RB_rv_superfoc7 PA=0

04:10 RB_rv_superfoc7_pa180

04:17 DOFPIA, /REDONLY

04:25 Turn off, turn on LBC because of consecutive shutter hardware failures. Issue 5287.

  • MODS image counter got a bit messed up. On red side, it says image 2 of 3 and on blue side it says image 3 of 3 but both say acquiring image 3 of 3... and we are on the third image.
  • One more spectrum.
  • UT 04:20, imaging mode, imcslock. Manually take through-slit images: 20-sec x2 (clear filters). 20-sec x 3 (g/i) filters.
    • collimation looking a bit odd on guider - jumping a lot.
    • 04:29 moving slit out and taking an exposure to check on-camera collimation. 20-sec x 3 at g and i. 2nd exposure is very elongated, first was ok. Another series of 3, offset -6" (detxy relative) after first one.
      • offset dY = -20 and repeat at g/i.
      • offset dY = +20 and u/r.
      • offset dX = +12 and u/r - collimation and seeing awful
  • Restarting symbiotic script, but shutter failures on every image. Stopped script, John going to cycle power on LBC.
    • 04:27 restarted symbiotic script. LBC-R shutter now failing every time.

04:55 Done with psuedo-monocular testing. Steve authorizing MODS on left and NONE on right.

05:00 execMODS NGC6341f1_PA180_ugrizPhot.img for check of photometric calibration on MODS1. Night is photometric.
  • airmass 1.284, seeing ~1.4-1.8"
  • Seeing improved to 1", so I repeated the script. Airmass now 1.32.
  • was trying to make a script for another Clem GC, NGC 6383, but it is a much more crowded field (see, e.g. 19:53:24.177 18:56:21.30 which is one position with many stars having u as well as griz... not all stars in Clem+2007 catalog have u photometry). In the end, decided to slew to spectrophot standard instead.

05:06 set Shutter to FALSE in redchannel.conf, then Turn off/on LBC.

Played OB, shutter worked 2 times before it failed.

05:19 lbckill/lbcstart

05:27 Turn off LBC; lbckill/lbcstart

05:27 acqMODS bd28... airmass 1.02.
modsAlign -r mods1r.20140902.0052.fits

MODS Offset Command:
   offsetxy -3.576  5.068 rel

Star not well centered in 53. Star appears at 502, but slit center seems to be 499 instead. dX = -0.4". Gave again +0.4", 55, but looks like 53. Gave -0.2" before starting dualgrating exposure.

05:30 Guide star images taken while collimating are weird -- star very elongated at first, then a donut. Seeing is 0.92".

05:57 bd28 done in dual grating.

05:57 Steve auth LBC-R on Right and MODS on Left.

05:59 acqMODS feige110.acq.
modsAlign -r mods1r.20140902.0059.fits
MODS Offset Command:
   offsetxy -3.576  5.972 rel
Again a negative dx offset required. Star is at X=502, but slit appears to be centered at X=499. Gave offset dx = -0.35 and took another thru-slit exposure, 61.

06:08 Turn on LBC

06:14 execMODS -b DualGrating ~modseng/modsScripts/modsSpecPhot/feige110.obs

06:23 in prep for observing a sci target for Barry, take feige110 through 0.6" slit also
  • acqMODS -i feige110.acq, actMODS -a feige110.acq (modified acq script in his directory)
  • 06:23: error. problem with tip-tilt request for left side
  • since we'd just centered the star already for the 5" slit and sent no preset for the 0.6" slit, I edited script to decrease exptime. The star is really already on the slit and modsAlign gave the following adjustment. Confirmatory image taken after the offset was applied is 69 - looked pretty good. But the mirror went down just now (06:35 UT).
MODS Offset Command:
   offsetxy  0.148 -0.109 rel

06:46 mirror raised. sending preset to fe110 again.
MODS Offset Command:
   offsetxy -3.872  5.489 rel
72 - needs shift of about 1 pix to the left. dx = -0.123 73, dx=-0.123 again. Next image, 74, looks good. So the total offset reqd was -0.246".

06:59 take a 300-sec spectrum on feige110 through the 0.6" slit. Seeing was about 0.8-0.9".

07:07 Steve authorizing MODS + none on the right side. LBC-Red shutter is not working again... earlier tonight it was intermittently failing, but then failed on every exposure.

07:10 acqMODS FF2131
MODS Offset Command:
   offsetxy -5.430  4.231 rel
Confirmatory image is 78, object slightly to right. Based on what I'd seen earlier for feige110, giving a detxy offset, dx = -0.2", take another thru-slit, 79. Looks a little better. Going to go with that. I centered on the object with "a", but if I'd eyeballed the center, the offset would have been:
-5.491  4.272 rel
, 0.06" different. And using the experimental version that does the bias-subtraction, I would get
 -5.496  4.246 rel
.

07:27 execMODS FF2131minus0141.obs
  • after 2 exposures, check slit centering by taking out slit, going to imaging mode and
    • go to red-only imaging mode at r-sdss with slit in.
      • 120-sec exposure: 82 guide star is looking like an "arrowhead"
      • 120-sec field image: 83 - objects are slightly elongated
    • no changes made to centering based on this pair of images, although on 83, the brightest spot may be to the left in the slit. Comparing field images 77 and 83, although stars on 83 appear elongated, the seeing is better, counts are higher, and there is more structure seen in the object itself.
  • again after 85, go to red-only imaging mode and take slit image (86), then field image (87).

09:22 acqMODS feige110 (will be at 1.4 airmass as it was before, will have to pick up the 1-1.4 and 1.6-2 airmasses at the end of the night).
MODS Offset Command:
   offsetxy -3.980  5.958 rel
star elongated but exposure 2-sec.

09:28 execMODS feige110 to obtain spectra in all modes. airmass 1.4-1.5.
  • guide star fwhm went up and then down around 10:00 UT, see SX guide plot.

10:18 acqMODS FF0139plus0115.acq
  • 104 looks well centered
MODS Offset Command:
   offsetxy -3.508  5.294 rel
  • execMODS FF0139plus0115.obs --- object has a nucleus which is much brighter than previous object, though still there is diffuse emission.
  • unfortunately, there isn't time to do the full observation. I cut the first one short at 10-min and will run a second 10-min exposure.

11:01 acqMODS feige110.acq (5" slit)
  • chose wrong guide star. There are two stars, one in each corner of the acq image. Steve correcting pointing.
MODS Offset Command:
   offsetxy -4.346  6.378 rel
  • confirmatory image, 109, looks good.

11:16 execMODS -b DualGrating feige110.obs (airmass 2.1)
  • red imcslock took longer than the timeout --- retry worked.
  • took 2 of the 3 exposures

11:28 acqMODS g191b2b.acq (just at 18 deg twilight now)
  • took a while to collimate
MODS Offset Command:
   offsetxy -3.010  5.767 rel
  • needed a dx offset, dx=-0.3" (115), and another, dx=-0.1" (116). Giving one more dx=-0.1" offset and running obs script.
  • airmass 1.140.
  • took a final thru-slit to confirm centering.

11:56 acqMODS gd71.acq
  • red imcslock imaging mode seems to be going slowly, but the counts are OK. Zero point was not changed much.
MODS Offset Command:
   offsetxy -3.066  5.644 rel
  • execMODS -b DualGrating gd71.obs
  • airmass 1.225
  • sky brightness now 17.59 however guiding is fine. Lot of twilight seen in spectrum --- will let this finish and that will be it.

Close @ 12:30

12:45 Dome dark, telescope at zenith. Taking some calibrations for the spectrophotometric standard done in all configurations.
  • grlamps, grlamps_r, grlamps_b
  • prlamps, prlamps_r, prlamps_b (via prlamps megascript)
    • Xe strongly saturated at the red end of blue
  • imflats_edit.cal:
    • imflats.cal script for MODS1 in ~modseng/modsScripts/modsCalib directory needs revision and I took some test exposures and ran this edited version (in ~lbto/SciPrep/MODS). Because of time, running only 1 image per filter.
filter vflat level exptime counts
u 10 50 21000
g 4 12 30000
r 3 2 27000
i 1.5 10 41000
z 1.5 4 39000

02:10 Calibrations done. Some work still needed at i and z.
  • Need prflats (dual, blue and red) and grpixflats (blue and red) still.

-- %USERSIG{JohnHill - 2014-09-02}%
Topic revision: r8 - 03 Sep 2014, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback